### v1

@@ 6.4.3 Mass flow rates
Finally, we considered the mass flows related to the reconstructed streaming motions. The distributions of $\dot{M}$, presented in Fig. 18, indicate a series of converging and diverging mass flows along the line of sight with standard deviations around 3.0 $\times$$10^{-3}$ and 0.6 $\times$$10^{-3}$ M/yr. It is beyond the scope of this work to determine whether there is a known MC complex at the interphase of all the Hi converging flows, as suggested by some of the MC formation scenarios (see, for example Koyama & Inutsuka, 2000). Reconstructing the specific flows in and around MC calls for dedicated studies of each region, as already done with the HOG method for the Taurus MC in (Soler et al., 2023). The map in Fig. 18 and the azimuthal profiles in Fig. 19 provide a framework to distinguish specific flows from the global fluctuations and establish a statistical comparison with the results of numerical simulation of the Galactic environment.
@@ 7 Conclusions
We presented a study of the morphological relation between a model of three-dimensional dust extinction and the Hi and CO line emission. We identified a significant correlation between dust and gas that we employed to study the ISM velocities across the region within $|b|$$<$ 5 and $d$$<$ 1250 pc. We produced maps of the LOS motions, streaming motions, kinetic energy and moment densities, and mass flow rate distributions across the studied volume.
We found that the material within 1250 pc from the Sun carries the global kinematic imprint of circular motions around the Galactic center. Yet, we also identified average departures from Galactic rotation, streaming motions, of around 12.1 and 6.1 km s-1 for the material sampled by Hi and CO, respectively.
@@
We found that the energy overdensities correspond to regions with momentum densities and mass flow rates above the mean. Yet it is difficult to assign these concentrations to individual feedback events or accumulations of material, as establishing those processes would require modeling the gas’s past configuration with the only velocity component available to observations. Yet the synergy of this kind of analysis with the reconstruction of the stellar dynamics enabled by Gaia and other observations and the novel techniques of data modeling promises to produce a complete picture of the ISM evolution in the Solar neighborhood and beyond.
@@ Acknowledgements.
JDS thanks the following people who helped with their encouragement and conversation: João Alves, Bruce Elmegreen, Torsten Eßlin, Álvaro Hacar, Sebastian Hutschenreuter, Claire Murray, Martin Piecka, and Eugenio Schisano. RSK also thanks the 2024/25 Class of Radcliffe Fellows for highly interesting and stimulating discussions. Some of the crucial discussions that led to this work took place under the program Milky-Way-Gaia of the PSI2 project, which is funded by the IDEX Paris-Saclay, ANR-11-IDEX-0003-02. This project was funded by the European Research Council via the ERC Synergy Grant “ECOGAL” (project ID 855130). RSK also acknowledges financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The computations for this work were performed at the Max-Planck Institute for Astronomy (MPIA) astro-node servers. RSK is also grateful for computing resources provided by the Ministry of Science, Research and the Arts (MWK) of the State of Baden-Württemberg through bwHPC and the German Science Foundation (DFG) through grants INST 35/1134-1 FUGG and 35/1597-1 FUGG, and also for data storage at SDS@hd funded through grants INST 35/1314-1 FUGG and INST 35/1503-1 FUGG.
Software: astropy (Astropy Collaboration et al., 2018), SciPy (Virtanen et al., 2020), astroHOG (Soler, 2020).
@@ A.1 Direction- and orientation-sensitive Rayleigh statistic
Fig. 25 compares $V$ and $V_{\rm d}$ for a test region. The distinction between parallel and antiparallel gradients increases the significance of the distance-$v_{\rm LOS}$ pairs with morphological correlation, remarkably increasing the contrast in comparing the 3D dust density and the CO emission. It also improves the comparison between the 3D dust density and the Hi emission by excluding outlying distance-$v_{\rm LOS}$ pairs that are antiparallel, as shown by their $V_{\rm d}$$<$00 values. These outliers may be produced by HiSA structures or chance correlation. The detailed study of these features requires further HiSA identification using spectral methods (see, for example, Gibson et al. 2005; Syed et al. 2023), which are beyond of the scope of this work.
@@ A.2.1 Line emission uncertainties
Figure 27 shows the $V_{\rm d}$ standard deviation for the example presented in Fig. 26. As expected from the negligible variations in the results obtained with $N_{\rm MC}$$=$ 0, 10, and 100, the values of $\sigma_{V_{\rm d}}$ are less than 1% below the mean values obtained with the Monte Carlo sampling. In essence, the maps in Fig. 27 reflect the significance of the emission detection across $v_{\rm LOS}$. Thus, the lower $\sigma_{V_{\rm d}}$ for the Hi observations are the result of the lower noise level in those observations, in comparison with the CO composite survey.
@@
The result of the logarithmic dust extinction modeling presented by Edenhofer et al. (2024) is a set of 12 samples from the variational posterior of the 3D dust distribution. We computed the correlation with the line emission data for each sample to propagate the reconstruction uncertainties into the HOG analysis. We reported the mean value of $V_{\rm d}$. This data model assumes that each sample is a plausible reconstruction of the 3D dust distribution and that the variance across the set represents the reconstruction’s uncertainties.
@@ A.3.3 Line-of-sight velocity range
In the main body of this paper, we chose the range $-25$$<$$v_{\rm LOS}$$<$ 25 km s-1 range for the input line emission data. This selection was motivated by the maximum LOS velocity expected from the Reid et al. (2019) Galactic rotation model within the heliocentric distances $d$$<$ 1.25 kpc, which is 21.6 km s-1. That assumption mitigates the effect of chance correlation in the morphological comparison, which due to the fixed angular resolution, is more acute for the largest distances and higher LOS velocities. This section considers the results obtained when this limit to $-40$$<$$v_{\rm LOS}$$<$ 40 km s-1.
Figure 32 shows the $v_{\rm LOS}$ reconstructions for the extended LOS velocity input range. The $v_{\rm LOS}$ maps preserve the quadrupolar pattern expected from the motions around the Galactic center, shown in Fig.  10. However, the reconstructed $v_{\rm LOS}$ amplitude is higher than in Fig. 32, implying larger departures from the purely rotational motions. These departures are illustrated in Fig. 33, where the differences between the reconstructed $v_{\rm LOS}$ and the expectation from the Reid et al. (2019) Galactic rotation model are reported.
The top panel in Fig. 33 shows that the distribution of streaming motions for the extended input range peaked is also around 0 kps. Still, there are many occurrences beyond the initial limit of $\pm$25 km s-1, particularly for Hi. These occurrences, however, are extended across multiple velocities, forming a wing pattern around the central distribution. We interpret this pattern as the effect of chance morphological correlation, which produces an accidental relation between the local 3D dust distribution and high-$v_{\rm LOS}$ line emission channels. This fact is further illustrated in the bottom panel of Fig. 33, where we report the results of the even larger input velocity range $-100$$<$$v_{\rm LOS}$$<$ 100 km s-1. In that case, the strong peak in streaming motions around 0 kps survives. Still, a significant signal is found at $v_{\rm LOS}$ beyond the range justifiable by the physical processes that could produce departures from Galactic rotation. Without more detailed studies, it is difficult to justify that these large departures are not the result of the limitations imposed by the angular resolution of the input data and the HOG method, so we conservatively reported the results for the narrower $v_{\rm LOS}$ range in the main body of the paper. Additional departures from circular rotation would produce larger velocity dispersion and, consequently, larger kinetic energy and momentum densities. Thus, our results should be considered lower limits for these quantities until more detailed reconstructions are produced.
@@
Figure 32: Same as Fig. 11, but for the input LOS velocity range $-40$$<$$v_{\rm LOS}$$<$ 40 km s-1.
Figure 33: Same as Fig. 12, but for the input LOS velocity ranges $-40$$<$$v_{\rm LOS}$$<$ 40 km s-1 and $-100$$<$$v_{\rm LOS}$$<$ 100 km s-1.
@@ Appendix B Hi self absorption and the HOG results
The plots in Fig. 5 and Fig. 6 present the maximum values of $V_{\rm d}$ identified across $v_{\rm LOS}$ channels for the volume cells in which we divided the studied region. This selection highlights structures where the density and line emission gradients are parallel, thus selecting regions with a positive correlation between emission intensity and density. However, there may be information in the minimum values of $V_{\rm d}$. If $V_{\rm d}$$<$$-2.87$ in a volume element, it may highlight a significant HISA feature, where the gradients for the dust density and Hi emission are antiparallel. To evaluate this possibility, we considered the values of $V_{\rm d}$ corresponding to the maximum $|V_{\rm d}|$ across distance channels, as we report in Fig. 34. That selection highlights distance channels dominated by antiparallel density and emission gradients.
Figure 35 shows an example of a distance channel with significantly antiparallel gradients, identified by $V_{\rm d}$$=$$-7.29$. Inspection of the Hi spectra toward the high-density areas in the right-hand-side panel of Fig. 35 indicates that the decrease in the Hi emission is not associated with a HISA feature. In this particular example and others throughout the studied volume, the minimum $V_{\rm d}$ values are not related to HISA but rather to the chance correlation between the Hi emission and the 3D dust. This means there is no physical significance in most of the minimum $V_{\rm d}$, which is why they are not considered in the main body of the text. The only way to systematically separate chance correlation from HISA is integrating the spectral information, which is not straightforward, as indicated in previous HISA studies (see, for example, Gibson et al. 2005; Syed et al. 2023), and is beyond the scope of this paper.
@@
Figure 35: Same as Fig. 4, but using $\max(|V_{\rm d}|)$ instead of $\max(V_{\rm d})$ for the assignment of $v_{\rm LOS}$ across distance channels.
Appendix C Calculation of physical quantities from the HOG method results
In the main body of this paper, we reported a set of physical quantities derived from the HOG analysis of the 3D dust and line emission data. The estimation of these quantities is based on a series of assumptions that we describe in detail in the following sections.
C.1 Dust density corresponding to Hi and CO emission
[…]
We found very few occurrences of $V_{\rm d}$$<$ 0 for the CO emission, indicating that the antiparallel emission and density gradients are rarely found. This is expected from the positive correlation between extinction and emission from that gas tracer at the optical depths sampled in the 3D dust reconstruction ($A_{\rm V}$$>$ 1) (see, for example, Glover & Mac Low 2011). The mean value of $V_{\rm d}$ for the CO data is below that significance level, but there is a considerable number of tiles with $V_{\rm d}$$>$ 2.87. This difference can be explained by the larger filling factor of the Hi emission in the same area of the sky, which means that there are more significant gradient vectors for Hi than for CO in the same tile. Yet, there are enough significant tiles in the comparison between 3D dust and CO emission to employ this gas tracer to reconstruct the motions within the volume covered in the Edenhofer et al. (2024) model.
Figure 36: Distribution of the $V_{\rm d}$ values presented in Fig. 5. The colored vertical lines correspond to the mean values for the corresponding line emission tracer.
The Edenhofer et al. (2024) 3D extinction model has a logarithmic binning of distances to account for the lower density of sources. This implies there are more tiles for the nearby channels than distant ones. To evaluate the potential bias that this feature may introduce in HOG results, we considered the distribution of $V_{\rm d}$ as a function of distance, illustrated in Fig. 37. Despite the expected clustering at lower distances, the values of $V_{\rm d}$ are not correlated with the distance to the tiles for either line emission tracer. Moreover, the most significant $V_{\rm d}$ values are not clustered around particular distances but are distributed along the line of sight, thus indicating that HOG is not biased by an effect related to the distance in the 3D dust model.

### v2

@@ 6.4.3 Mass flow rates
Finally, we considered the mass flows related to the reconstructed streaming motions. The distributions of $\dot{M}$, presented in Fig. 18, indicate a series of converging and diverging mass flows along the line of sight with standard deviations around 3.0 $\times$$10^{-3}$ and 1.2 $\times$$10^{-3}$ M/yr. It is beyond the scope of this work to determine whether there is a known MC complex at the interphase of all the Hi converging flows, as suggested by some of the MC formation scenarios (see, for example, Koyama & Inutsuka, 2000). Reconstructing the specific flows in and around MC calls for dedicated studies of each region, as done with the HOG method for the Taurus MC in Soler et al. (2023). The map in Fig. 18 and the azimuthal profiles in Fig. 19 provide a framework to distinguish specific flows from the global fluctuations and establish a statistical comparison with the results of numerical simulation of the Galactic environment.
[…]
@@ 7 Conclusions
In this paper, we studied the morphological relation between a model of three-dimensional dust extinction and the Hi and CO line emission. We identified a significant correlation between dust and gas that we employed to study the ISM velocities across the region within $|b|$$<$ 5 and $d$$<$ 1250 pc. We produced maps of the LOS motions, streaming motions, kinetic energy and moment densities, and mass flow rate distributions across the studied volume.
We found that the material within 1250 pc from the Sun carries the global kinematic imprint of circular motions around the Galactic center. Yet, we also identified average departures from Galactic rotation, streaming motions, of around 10.8 and 6.6 km s-1 for the material sampled by Hi and CO, respectively.
@@
We found that the energy overdensities correspond to regions with momentum densities and mass flow rates above the mean. It is difficult to assign these concentrations to individual feedback events or accumulations of material, as establishing those processes would require modeling the gas’s past configuration with the only velocity component available to observations. However, the synergy of this kind of analysis with the reconstruction of the stellar dynamics enabled by Gaia and other observations and the novel techniques of data modeling promises to produce a complete picture of the ISM evolution in the Solar neighborhood and beyond.
@@ Acknowledgements.
We thank the anonymous referee for carefully reading our manuscript and providing insightful comments and suggestions. JDS thanks the following people who helped with their encouragement and conversation: João Alves, Bruce Elmegreen, Torsten Eßlin, Álvaro Hacar, Sebastian Hutschenreuter, Claire Murray, Martin Piecka, Eugenio Schisano, Mark Heyer, and Oleg Egorov. JDS also thanks Daniel Seifried and Stefanie Walch-Gassner for providing the numerical simulations and synthetic observations to test the HOG method. RSK also thanks the 2024/25 Class of Radcliffe Fellows for the stimulating discussions. Some of the crucial discussions that led to this work took place under the program Milky-Way-Gaia of the PSI2 project, which is funded by the IDEX Paris-Saclay, ANR-11-IDEX-0003-02. This project was funded by the European Research Council via the ERC Synergy Grant “ECOGAL” (project ID 855130). RSK also acknowledges financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The computations for this work were performed at the Max-Planck Institute for Astronomy (MPIA) astro-node servers. RSK is also grateful for computing resources provided by the Ministry of Science, Research and the Arts (MWK) of the State of Baden-Württemberg through bwHPC and the German Science Foundation (DFG) through grants INST 35/1134-1 FUGG and 35/1597-1 FUGG, and also for data storage at SDS@hd funded through grants INST 35/1314-1 FUGG and INST 35/1503-1 FUGG. Software: astropy (Astropy Collaboration et al., 2018), SciPy (Virtanen et al., 2020), astroHOG (Soler, 2020).
@@ A.1 Direction- and orientation-sensitive Rayleigh statistic
Fig. 25 compares $V$ and $V_{\rm d}$ for a test region. The distinction between parallel and antiparallel gradients increases the significance of the distance-$v_{\rm LOS}$ pairs with morphological correlation, remarkably increasing the contrast in comparing the 3D dust density and the CO emission. It also improves the comparison between the 3D dust density and the Hi emission by excluding outlying distance-$v_{\rm LOS}$ pairs that are antiparallel, as shown by their $V_{\rm d}$$<$00 values. These outliers may be produced by HiSA structures or chance correlation. The detailed study of these features requires further HiSA identification using spectral methods (see, for example, Gibson et al. 2005; Syed et al. 2023), which are beyond the scope of this work.
@@ A.2.1 Line emission uncertainties
Figure 27 presents the $V_{\rm d}$ standard deviation for the example presented in Fig. 26. As expected from the negligible variations in the results obtained with $N_{\rm MC}$$=$ 0, 10, and 100, the values of $\sigma_{V_{\rm d}}$ are less than 1% below the mean values obtained with the Monte Carlo sampling. In essence, the maps in Fig. 27 reflect the significance of the emission detection across $v_{\rm LOS}$. Thus, the lower $\sigma_{V_{\rm d}}$ for the Hi observations are the result of the lower noise level in those observations, in comparison with the CO composite survey.
@@
The dust extinction modeling result presented in Edenhofer et al. (2024) is a set of 12 samples from the variational posterior of the 3D dust distribution. We computed the correlation with the line emission data for each sample to propagate the reconstruction uncertainties into the HOG analysis. We reported the mean value of $V_{\rm d}$. This data model assumes that each sample is a plausible reconstruction of the 3D dust distribution and that the variance across the set represents the reconstruction’s uncertainties.
@@
A.2.3 Chance correlation
A significant effect to consider in the morphological matching using the HOG method is the chance correlation, which is the increase in $V_{\rm d}$ due to the accidental similarity between images. For example, most landscape photographs taken by humans on the surface of the Earth are correlated by chance due to the presence of the horizon. In the same way, the vertical symmetry of the Galactic plane imposes a general level of morphological correlation in all tracers. This effect can be minimized by employing the highest possible number of independent gradient vectors to resolve individual characteristics in the two images and produce contrast above the chance correlation level. In our application, the number of independent gradient vectors is fixed by the 3D dust reconstruction’s angular resolution. Thus, we employ additional tests to marginalize the effect of change correlation.
Figures 30 and 31 present the $V_{\rm d}$ values obtained for the test tile in Fig. 4 when flipping the emission PPV cubes with respect to the 3D dust PPD cube in the vertical (JK01), horizontal (JK10), and diagonal (JK11) directions. The JK label comes from Soler et al. (2019), where this operation was introduced as “jackknife” tests. The flipped cubes have the same global statistical properties, but their morphology should differ from the originals. That difference level is quantified by the $V_{d}$ values obtained with the flipped sets. For example, if the line emission is distributed like a large symmetric blob, it would produce a similar $V_{d}$ in the three flip tests, thus suggesting that it has insufficient unique features to produce an unambiguous morphological match.
The root-mean-square of the jackknife tests, $\sigma^{\rm JK}_{V_{\rm d}}$, reveals a significant number of $d$-$v_{\rm LOS}$ pairs where the chance correlation can be acute, as illustrated in Fig. 30 and Fig. 31. This is expected, given the multiscale correlations in the ISM that can mimic similarity when observed or reconstructed at low angular resolution. We account for this effect in $v_{\rm LOS}$ reconstruction by evaluating the analysis results against the jackknife tests’ amplitude, as described in Sec. 3.
Figure 31: Same as Fig. 30, but for CO.
@@ A.3.2 Galactic plane segmentation
We evaluated the effect of the Galactic plane segmentation in the $v_{\rm LOS}$ assignment by considering the results obtained for three $l$ divisions. Figure 34 presents the distribution of the $v_{\rm LOS}$ assigned to the dust tiles using the Hi and CO emission. We found that the reconstructed $v_{\rm LOS}$ have a global distribution primarily unaffected by the sky segmentation. The discrepancies between the three $v_{\rm LOS}$ distributions indicate the suppression of some velocity features toward particular lines of sight, which are worth exploring in a focused study of specific regions but do not significantly change the conclusions of our work.
@@ A.3.3 Line-of-sight velocity range
In the main body of this paper, we chose the range $-25$$<$$v_{\rm LOS}$$<$ 25 km s-1 for the line emission data input. This selection was initially motivated by the maximum LOS velocity expected from the Reid et al. (2019) Galactic rotation model within the heliocentric distances $d$$<$ 1.25 kpc, which is 21.6 km s-1. This restricted input $v_{\rm LOS}$ range aims to mitigate the effect of chance correlation in the morphological comparison since, in principle, line emission at higher $v_{\rm LOS}$ is not necessarily related to the local material. This section tests that assumption by considering the results obtained with broader $v_{\rm LOS}$ input ranges..
Figure 35 presents the distribution of the maximum $V_{\rm d}$ for the comparison between the 3D dust and the Hi and CO line emission in the range $-80$$<$$v_{\rm LOS}$$<$ 80 km s-1. Visual comparison between Fig. 5 and Fig. 35 demonstrates that the extension of the $v_{\rm LOS}$ input range does not increase the morphological correlation, quantified by $\max(V_{\rm d})_{V_{\rm LOS}}$. This result suggests that the low-correlation regions, $\max(V_{\rm d})_{V_{\rm LOS}}$$\approx$ 0, are not the consequences of the limited $v_{\rm LOS}$-input but a result of the 3D dust distribution across distance channels.
@@
Figure 35: Same as Fig. 5, but for the input LOS velocity ranges $-120$$<$$v_{\rm LOS}$$<$ 120 km s-1.
Figure 36 shows the results of expanded LOS velocity input ranges in the distance-$v_{\rm LOS}$ diagrams introduced in Fig. 4. For the region in that example, a significant morphological correlation exists for Hi beyond the $|v_{\rm LOS}|$$<$ 25 km s-1limit. Except for a few outliers, the most significant morphological correlation for CO is within the range $|v_{\rm LOS}|$$<$ 25 km s-1. This result implies that the $v_{\rm LOS}$ input limit excludes some significant departures from circular motions, effectively limiting the amplitude of the reconstructed streaming motions and the energy and momentum estimates. However, the HOG method also indicates contamination by chance correlation when employing a broader $v_{\rm LOS}$ range.
Figure 36: Same as Fig. 4, but for the input LOS velocity range $-120$$<$$v_{\rm LOS}$$<$ 120 km s-1.
Figure 37 presents the distribution of the streaming motions derived with the extended input range $-120$$<$$v_{\rm LOS}$$<$ 120 km s-1. The histogram shows a background signal in Hi that extends beyond the range $|v_{\rm LOS}|$$<$ 25 km s-1, particularly for negative $v_{\rm LOS}$. This signal is unlikely to have a natural origin, as it would imply a significant velocity gradient in the solar vicinity, in tension with the observations of other tracers and the conditions for the Galactic disk stability. It is more plausible that this signal results from accidental correlation remanent after the selection described in App. A.2.3. The CO emission appears less affected by this additional chance correlation, so we use it as a reference to set the input $v_{\rm LOS}$ limits of the reconstruction. For the main body of the paper, we chose an input $v_{\rm LOS}$ range centered on zero and spanning across three times the standard deviation of the CO streaming motions estimated with the extended for the range $-120$$<$$v_{\rm LOS}$$<$$120$ km s-1, which is 52.1 km s-1. Hence the resulting $|v_{\rm LOS}|$$<$ 25 km s-1 range. We reserve the study of broader departures from circular motion to a subsequent, less general study.
Figure 37: Same as Fig. 12, but for the input LOS velocity range $-120$$<$$v_{\rm LOS}$$<$ 120 km s-1.
@@ Appendix B Hi self absorption and the HOG results
The plots in Fig. 5 and Fig. 6 present the maximum values of $V_{\rm d}$ identified across $v_{\rm LOS}$ channels for the volume cells in which we divided the studied region. This selection highlights structures where the density and line emission gradients are parallel, thus highlighting a positive correlation between emission intensity and density. However, there may be information in the minimum values of $V_{\rm d}$, which in the case of significant negative values may highlight a HISA feature, where the gradients for the dust density and Hi emission are antiparallel. To evaluate this possibility, we considered the values of $V_{\rm d}$ corresponding to the maximum $|V_{\rm d}|$ across distance channels, as we report in Fig. 38. That selection accentuates distance channels dominated by antiparallel density and emission gradients.
@@
Figure 38 presents the $V_{\rm d}$ values corresponding to the maximum $|V_{\rm d}|$ across distance channels in the studied volume. Visual comparison with Fig 5 shows that, in general, the negative $V_{\rm d}$ are only dominant in regions with relatively low $\max(V_{\rm d})$. That result implies the positive correlation between 3D dust density and Hi emission dominates throughout the studied volume. However, we performed additional tests to determine whether dust structures are associated with HISA in the distance channels with significant $V_{\rm d}$$<$ 0.
Figure 39: Same as Fig. 4, but also indicating the maximum $|V_{\rm d}|$ for Hi in each distance channel (magenta markers). The dashed vertical and horizontal lines indicate the distance and $v_{\rm LOS}$ with the minimum $V_{\rm d}$ toward this region.
Figure 39 show the morphological correlation between 3D dust distance channels and line-emission $v_{\rm LOS}$ channels, indicating the $v_{\rm LOS}$ with the maximum $|V_{\rm d}|$ at each distance after the selection criteria described in Sec. 3. As expected from the results in Fig. 38, in most distance channels, the maximum $|V_{\rm d}|$ coincides with the maximum $V_{\rm d}$. However, in this example, the distance channel with the minimum $V_{\rm d}$ corresponds to a $v_{\rm LOS}$ that would have been excluded in selecting positive correlation by maximum $V_{\rm d}$.
The top panel of Fig. 40 presents the line emission and 3D dust density for the channels with the minimum $V_{\rm d}$ identified in Fig. 39. It is apparent from the images that there is a decrease in the Hi intensity toward the region with high CO emission on the central-left part of the maps. Inspection of the Hi spectrum confirms the presence of a few-km s-1-wide dip that usually characterizes HISA (Gibson et al. 2000). A large fraction of the $V_{\rm d}$ signal comes from the morphological coincidence between the depression in the Hi emission and the 3D structures toward the lower portion of the map. The spectra toward that portion of the map are not straightforward to interpret but show some promising indications of HISA. However, a generalization of this result is complicated.
The bottom panel of Fig. 40 displays the line emission and 3D dust density for the distance-$v_{\rm LOS}$ pair with the lowest $V_{\rm d}$ in the studied volume. The gradient in 3D dust density coincides with the Hi emission decrease. In this case, however, the depression in the Hi emission is extended tens of km s-1, thus it is unlikely to come from HISA. Our results can be used to indicate the potential HISA locations, but their confirmation requires additional spectral analysis beyond the scope of this work. Consequently, we report only the results for the positive correlation between Hi emission and 3D dust in the main body of the paper.
Figure 40: Two examples of regions where the Hi emission gradients are mainly antiparallel to the 3D dust density gradients, as identified by the minimum values of $V_{d}$.
Appendix C Effects of fixed angular resolution and distance
In this section, we consider the effect of the fixed angular resolution introduced by the derivative kernel size in the $V_{\rm d}$ distribution and its potential effect in the results of the HOG method. Our selection of a 0.5 derivative kernel implies that we are sampling scales around 0.6 pc at the front of the 3D dust reconstruction, $d$$=$ 69 pc, and 10.9 pc at the back, $d$$=$ 1250 pc. We evaluate the potential effect of this feature in the reported $V_{\rm d}$ values by considering its distribution across distance channels.
Figure 41 presents the distribution of maximum $V_{\rm d}$ values reported in Fig. 5, but divided in distance bins. The trends obtained for the 50th and 99th percentile of this quantity indicate that the fixed angular resolution does not produce a systematic fall in the $\max(V_{\rm d})$ values, but that these rather increase with distance. This increasing trend is most likely the result of the inclusion of additional ISM structures in the most distant channels compensating for the lack of spatial resolution. These results are not unexpected, considering the self-similar nature of the ISM, which implies that the correlation between Hi, CO, and 3D dust is extended across the scales considered in this study. Figure 41 indicates that the highest $\max(V_{\rm d})$ values are not clustered around particular distances but are distributed along the line of sight, thus indicating that HOG is not biased by an effect related to the distance in the 3D dust model.